The value of a certain two digit number is 9 times the sum of it's digits. If the digits are reversed, the resulting number is 63 less than the original number. What is the original number? Please, please help me answer this question!
Found 2 solutions by fcabanski, stanbon:
Answer by fcabanski(1391)

Look at the second part first. xy - 63 = yx. Look at the possibilities for y-3=x:


y=1 so 1-3=x=-2 or, if a ten had been carried from the tens digit it's 11-3=8.


Could x=8? 8, carry one ten to the ones digit so that becomes 7-6 = 1.


Yes, if x=8 and y=1 it's 81-63=18.


Check it in the first part.


xy = 81 = 9 (8+1)=9*9. Yes!

If you need help understanding math so you can solve these problems
yourself, then one on one online tutoring is the answer ($30/hr). If you
need faster solutions with guaranteed detailed answers, then go with
personal problem solving ($3.50-$5.50 per problem). Contact me at
fcabanski@hotmail.com


Answer by stanbon(75887)

The value of a certain two digit number is 9 times the sum of it's digits. If the digits are reversed, the resulting number is 63 less than the original number.
----
Let the number be 10t+u.
----
Equations:
10t+u = 9(t+u)
10u+t = 10t+u-63
-----
Simplify each equation:
t = 8u
9t = 9u + 63
-----
Modify:
t = 8u
t = u + 7
---
Solve for "u":
8u = u+7
u = 1
----
Solve for "t":
t = 8u
t = 8
---
Original Number: 10t+u = 81
=================================
Cheers,
Stan H.
========================